the given points are: A(1,3) and B(4,-1) whats the equation of ab?
Answer by stanbon(75887)

A(1,3) and B(4,-1) what's the equation of ab
---------------------
slope = (-1-3)/(4-1) = -4/3
--------------------------------
Form of the equation: y = mx + b
Since y= 3 when x = 1 and since m=-4/3 you can solve for "b":
3 = (-4/3)*1 + b
b = 13/3
-----------
Equation: y = (-4/3)x + (13/3)
==========================
Cheers,
Stan H.